OpenStudy (anonymous):

If a and b are positive integers and a2 = 49 and 3b3 = 192. Find the value of 2a + 3b.

OpenStudy (anonymous):

is it 26

OpenStudy (anonymous):

Is this your question \[\large a^2=49 \space \space and \space \space 3b^3=192\] ??

OpenStudy (anonymous):

if this is ur question in that case \[a=7 , b = 4\]

OpenStudy (anonymous):

\[2a+3b=2\times{7} + 3\times{4}=26\]

OpenStudy (anonymous):

TRUE